# arXiv:2603.01290v1[cs.AI] 01 Mar 2026

### v1

@@ 4.1 Conditional Independence: A Tractable Baseline
We adopt conditional independence as a deliberate tractable baseline — not a minor concession, but a principled starting point whose limitations are precisely characterised here and whose correct fix (a bivariate Gaussian over $\Delta v_{\text{trap}}$ and $\Delta t_{\text{sector}}$) is identified and deferred to Paper 2’s richer emission model:
$P(o_{t}\mid x_{t},z_{\text{zone}})=\prod_{i=1}^{5}P(o_{t,i}\mid x_{t},z_{\text{zone}})$ (6)
$\Delta v_{\text{trap}}$ and $\Delta t_{\text{sector}}$ are mechanically coupled: deploying more ERS simultaneously increases speed and decreases sector time, so the two observables carry correlated evidence about ERS state. $\Delta b_{\text{brake}}$ is also mechanically coupled to both. Treating them as conditionally independent double-counts this correlated evidence on every belief update, producing systematically over-confident posteriors.
The ECE =0.11=0.11 figure reported in Section 7 is measured on synthetic data where the independence assumption holds by construction — the synthetic generative model also assumes independence. This means 0.11 is a lower bound on the real-data calibration error, not an estimate of it. On real telemetry, where the mechanical coupling between observables is genuine, ECE is expected to be substantially higher. Baum-Welch EM on real sequences will absorb some of this correlation into the emission parameters, but cannot fully correct for a structural independence violation. Temperature scaling (Guo et al., 2017) is available as a post-hoc calibration step after Melbourne data is available.
@@ 4.2 Observable Distributions
Observables 1–4 are modelled as Gaussian, observable 5 as Bernoulli. The emission matrix is computed analytically by integrating these distributions over discretisation bin boundaries (Section 4.4).
Observable 1: $\Delta v_{\text{trap}}$ (km/h).
Primary drivers: ERS level and Active Aero deployment.
Table 1: $\Delta v_{\text{trap}}$ emission parameters. Non-activation zones.
ERS MOM $\mu$ $\sigma$ Interpretation
@@
L any 2.0-2.0 1.5 Harvesting
MOM active (transient event): $\mu=+7.0$, $\sigma=1.0$ (Sharp power spike)
@@
Circuit caveat.
The emission parameters above assume that $\Delta v_{\text{trap}}<0$ is an informative signal of below-baseline ERS deployment. This holds at high-regen circuits. At low-regen circuits where the recharge-per-lap ratio is $\leq 1.0$ (Australia, Italy, Saudi Arabia), the 5-lap rolling baseline shifts downward as universal super clipping becomes the modal operating mode, and the L-state emission mean of 2.0-2.0 km/h may underestimate the magnitude of the speed reduction during mandatory super-clipping phases. Baum-Welch recalibration is expected to absorb this circuit-specific shift into the emission parameters from Race 1 onwards.
Observable 2: $\Delta t_{\text{sector}}$ (seconds).
Primary drivers: ERS level and tyre degradation grade.
$\mu$ $=\mu_{\text{ERS}}+\delta_{\tau}\cdot f_{\text{ers}}$ (7)
$\mu_{\text{ERS}}$ $\in\{-0.30,0.00,+0.40\}\text{ s for ERS}\in\{H,M,L\}$ (8)
$\delta_{\tau}$ $\in\{0.0,0.1,0.25,0.50,0.90\}\text{ s for tyre}\in\{\text{new, light, moderate, heavy, cliff}\}$ (9)
$f_{\text{ers}}$ $\in\{0.6,1.0,1.3\}\text{ for ERS}\in\{H,M,L\}$ (10)
@@
Observable 3: $\Delta b_{\text{brake}}$ (metres).
Primary drivers: ERS regeneration braking point and tyre grip degradation. ERS regen advances braking distance by $\{+8,0,-10\}$m for H/M/L. Tyre degradation increases late-braking requirement by $\{0,3,8,15,24\}$m. Standard deviation $\sigma_{\tau}\in\{8,8,9,10,12\}$m.
Observable 4: $\sigma^{2}_{\text{speed}}$ (detrended variance).
Primary driver: ERS level. $\mu\in\{0.05,0.10,0.20\}$ for H/M/L. Override spent state contributes a +0.02+0.02 adjustment. MOM active: $\mu=0.03$, $\sigma=0.01$.
Observable 5: $z_{\text{aero}}$ (Bernoulli).
$P(z_{\text{aero}}=1\mid z_{\text{zone}}=0)=0$. Inside activation zones: $P(z_{\text{aero}}=1\mid z_{\text{zone}}=1)\in\{0.85,0.70,0.80\}$ for ERS $\{H,M,L\}$.
@@ 4.3 Emission Signal Strength: Audit Results
We audited the emission model to characterise which components of the hidden state are genuinely identifiable from observations.
ERS is strongly identifiable: $|\mu_{H}-\mu_{L}|=6.0$ km/h in $\Delta v_{\text{trap}}$ and 18m in $\Delta b_{\text{brake}}$, both substantially larger than the emission noise.
Tyre degradation signal is concentrated in the cliff state. Under normal degradation (new $\to$ heavy), the rolling 5-lap baseline absorbs the smooth monotone pace evolution. Cliff produces a step change the baseline cannot absorb ($\Delta t_{\text{sector}}\approx+0.8$s, $\Delta b_{\text{brake}}\approx+16$m, both clearly sub-baseline).
Override Mode status (available / spent) is not directly distinguishable from emissions in the non-active state. The differences in $\Delta v_{\text{trap}}$ between states are <0.5<0.5 km/h. The Override activation event is strongly identifiable ($\Delta v_{\text{trap}}\approx+7$ km/h, $\Delta t_{\text{sector}}\approx-0.65$s).
@@ 4.4 Observation Discretisation
Observations are discretised to form a finite joint observation space: $\Delta v_{\text{trap}}$: 8 bins; $\Delta t_{\text{sector}}$: 7 bins; $\Delta b_{\text{brake}}$: 5 bins; $\sigma^{2}_{\text{speed}}$: 5 bins; $z_{\text{aero}}$: 2 bins. Combined space: $8\times 7\times 5\times 5\times 2=2{,}800$ bins. Full bin boundaries are given in Appendix A.
The emission matrix $E\in\mathbb{R}^{30\times 2800}$ is computed analytically by integrating the Gaussian and Bernoulli distributions over bin boundaries, conditioning on each of the 30 hidden states.
@@ 5 Inference Architecture
The inference system is a two-layer pipeline trained in two phases and updated on an alternating schedule tied to the race calendar.
@@ 5.1 Layer 1: The HMM Belief Update
The HMM has 30 hidden states, emission matrix $E\in\mathbb{R}^{30\times 2800}$, transition matrix $T\in\mathbb{R}^{30\times 30}$, and initial distribution $\pi\in\Delta^{30}$.
@@
The transition probability $T_{ij}=P(x_{t+1}=j\mid x_{t}=i)$ factors over the three components under conditional independence:
@@
The ERS transition is marginalised over the rival’s action with estimated $P(\text{burn})$:
$P(e^{\prime}\mid e)=p_{\text{burn}}\cdot T^{\text{burn}}_{ee^{\prime}}+(1-p_{\text{burn}})\cdot T^{\text{harvest}}_{ee^{\prime}}$ (12)
The Override Mode transition is proximity-conditioned. $P(m^{\prime}=\text{available}\mid m=\text{spent})=P(\text{gap}<1\text{s})$, which is partially observable from live timing data. The tyre transition is monotone-absorbing. Full transition matrices are given in Appendix B.
@@
Phase 1 (pre-Melbourne): $E$ is set analytically; $T$ from physics-derived transition parameters; $\pi$ concentrates on $(e=H,m=\text{available},\tau=\text{new})$. Phase 2 (post-Melbourne): Baum-Welch EM (Baum et al., 1970) updates $E$ and $T$ from real race sequences.
@@ 5.2 Layer 2: The POMDP Policy
Input vector (55 dimensions).
Table 2: DQN input vector components.
@@
$e_{\text{self}}$ one-hot 3 Own ERS level
$m_{\text{self}}$ one-hot 2 Own Override Mode status
$f_{\text{self}}$ one-hot 3 Fuel (from lap number)
gap_ahead one-hot 4 close / threat / neutral / clear
gap_behind one-hot 4
[…]
tyre_compound one-hot 3 soft / medium / hard
tyre_age scalar 1 Normalised 0–1
lap scalar 1 Normalised 0–1
sector one-hot 3
$b_{t}$ probability vector 30 HMM belief state
Total 55
@@
Input(55) $\to$ Dense(256, ReLU) $\to$ Dense(256, ReLU) $\to$ Dense(128, ReLU) $\to$ Dense(2)
@@
Double DQN (van Hasselt et al., 2016) with experience replay (buffer capacity 50,000; Huber loss; gradient clipping at 10). $\epsilon$-greedy exploration decays from 1.0 to 0.05 over synthetic pre-training.
@@
During synthetic pre-training only:
$R_{\text{shaped}}$ $=R+\lambda\cdot\Phi,\quad\lambda=0.3$ (14)
$\Phi$ $=+0.10\cdot\mathbf{1}[\text{MAP}(b_{t})=x^{*}_{t}]$
$\quad-0.20\cdot\mathbf{1}[\text{trap condition}\wedge a=\text{burn}]$
$\quad+0.05\cdot\mathbf{1}[\text{gap\_ahead}\in\{\text{close, threat}\}\wedge e_{\text{self}}\geq M]$ (15)
@@
The DQN receives the HMM belief state $b_{t}$ rather than raw observables ($\Delta v_{\text{trap}}$, $z_{\text{aero}}$). This is theoretically correct: in a POMDP, the belief state is a sufficient statistic of the entire observation history. Given correct HMM parameters, the policy loses no information by operating on $b_{t}$ rather than the raw signal; in practice, however, an incorrect or surprised HMM produces a biased $b_{t}$, and the DQN then acts on that bias with no direct corrective — specifically, the residuals the HMM could not explain are not visible to the policy. End-to-end learning directly from raw observables to actions was considered but rejected in favour of POMDP coherence and interpretability: passing raw observables alongside $b_{t}$ would double-count evidence already integrated into the belief state, violating the sufficient-statistic property. Two partial mitigations are in place: (1) the $\delta\text{gap}$ input provides a raw, HMM-independent signal for closure rate; and (2) trap detection is explicitly computed as a function of $b_{t}$ and the zone flag, not a raw-signal threshold. A richer architecture that handles HMM residuals explicitly is deferred to Paper 2.
@@ 6.1 Data Source
Race telemetry is retrieved via the FastF1 Python library (Schaefer, 2020), which provides access to the official FOM timing and telemetry feed at 10 Hz sampling. The car_data endpoint provides per-car speed, throttle, and braking data at sector resolution, available for race sessions from 2018 onwards.

### v2

@@ 4.1 Conditional Independence: A Tractable Baseline
We adopt conditional independence as a principled tractable baseline:
$P(o_{t}\mid x_{t},z_{\mathrm{zone}})=\prod_{i=1}^{6}P(o_{t,i}\mid x_{t},z_{\mathrm{zone}})$ (7)
$\Delta v_{\mathrm{trap}}$ and $\Delta t_{\mathrm{sector}}$ are mechanically coupled: deploying more ers simultaneously increases speed and decreases sector time. Treating them as conditionally independent double-counts this correlated evidence on every belief update, producing systematically over-confident posteriors. The ECE figures reported in Section 7 are measured on synthetic data where the independence assumption holds by construction—they are lower bounds on real-data calibration error. Temperature scaling (Guo et al., 2017) is available as a post-hoc calibration step after Melbourne data is available.
@@ 4.2 Observable Distributions
Observables 1–4 are modelled as Gaussian, observable 5 as Bernoulli, and observable 6 as a truncated Gaussian on $[0,1]$.
Observable 1: $\Delta v_{\mathrm{trap}}$ (km/h).
Table 1: $\Delta v_{\mathrm{trap}}$ emission parameters. Non-activation zones.
ERS mode MOM $\mu$ $\sigma$ Interpretation
@@
$L_{\mathrm{harvest}}$ any 2.0-2.0 $1.5$ Deliberate harvest; managed throttle
$L_{\mathrm{derate}}$ any 1.5-1.5 $1.5$ SOC ceiling; slightly less depressed than harvest
MOM active (transient): $\mu=+7.0$, $\sigma=1.0$
@@
Circuit caveat. At low-regen circuits (recharge ratio $\leq 1.0$), the rolling baseline shifts downward as $L_{\mathrm{derate}}$ becomes the modal operating mode. Baum-Welch recalibration is expected to absorb this circuit-specific shift from Race 1 onwards.
Observable 2: $\Delta t_{\mathrm{sector}}$ (seconds).
$\mu$ $=\mu_{\textsc{ers}}+\delta_{\tau}\cdot f_{\textsc{ers}}$ (8)
$\mu_{\textsc{ers}}$ $\in\{-0.30,\;0.00,\;+0.40,\;+0.30\}\,\text{s}\quad\text{for }\textsc{ers}\in\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$ (9)
$\delta_{\tau}$ $\in\{0.0,\;0.1,\;0.25,\;0.50,\;0.90\}\,\text{s}\quad\text{for tyre}\in\{\mathtt{new},\mathtt{light},\mathtt{moderate},\mathtt{heavy},\mathtt{cliff}\}$ (10)
$f_{\textsc{ers}}$ $\in\{0.6,\;1.0,\;1.3,\;1.2\}\quad\text{for }\textsc{ers}\in\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$ (11)
@@
The $\mu_{\textsc{ers}}$ for $L_{\mathrm{derate}}$ (+0.30+0.30 s) is slightly less than for $L_{\mathrm{harvest}}$ (+0.40+0.40 s) because the car in $L_{\mathrm{derate}}$ is running full throttle; the time penalty comes from missing electrical power, not from deliberate harvesting.
Observable 3: $\Delta b_{\mathrm{brake}}$ (metres).
Primary drivers: ers regeneration braking point and tyre grip degradation. ers regen advances braking distance by $\{+8,0,-10,-5\}$ m for $\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$. The $L_{\mathrm{derate}}$ value of 5-5 m reflects that the car brakes marginally earlier under derating (less engine braking assistance at corner entry). Tyre degradation increases late-braking requirement by $\{0,3,8,15,24\}$ m. Standard deviation $\sigma_{\tau}\in\{8,8,9,10,12\}$ m.
Observable 4: $\sigma^{2}_{\mathrm{speed}}$ (detrended variance).
Primary driver: ers level. $\mu\in\{0.05,0.10,0.20,0.22\}$ for $\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$. The $L_{\mathrm{derate}}$ value is slightly elevated over $L_{\mathrm{harvest}}$ because intermittent SOC-ceiling events produce speed trace irregularities. Override spent state contributes a +0.02+0.02 adjustment. MOM active: $\mu=0.03$, $\sigma=0.01$.
Observable 5: $z_{\mathrm{aero}}$ (Bernoulli).
$P(z_{\mathrm{aero}}=1\mid z_{\mathrm{zone}}=0)=0$. Inside activation zones: $P(z_{\mathrm{aero}}=1\mid z_{\mathrm{zone}}=1)\in\{0.85,0.70,0.80,0.82\}$ for $\textsc{ers}\in\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$. The probability for $L_{\mathrm{derate}}$ is slightly higher than $L_{\mathrm{harvest}}$ because a derating car maintains full throttle—which is compatible with Active Aero deployment—whereas a harvesting car may modulate aero strategy with throttle management.
Observable 6: $\delta_{\mathrm{throttle}}$ (super-clipping duration fraction).
This is the primary discriminator between $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$, and the architectural motivation for the 40-state decomposition:
$\delta_{\mathrm{throttle}}=\frac{1}{|S|}\sum_{k\in S}\mathbf{1}\!\left[\mathrm{throttle}_{k}\geq 0.98\;\wedge\;v_{k}<\bar{v}\right]$ (12)
where $S$ is the set of straight-line telemetry samples in the sector, $v_{k}$ is instantaneous speed, and $\bar{v}$ is the driver’s 5-lap rolling speed baseline for that sector.
Table 2: $\delta_{\mathrm{throttle}}$ emission parameters. Values are Gaussian $(\mu,\sigma)$ truncated to $[0,1]$.
ERS state Sub-mode $\mu$ $\sigma$ Interpretation
$H$ $0.05$ $0.04$ Battery not depleted; rarely clips
$M$ $0.15$ $0.08$ Occasional clipping on long straights
$L_{\mathrm{harvest}}$ deliberate $0.08$ $0.05$ Managed throttle; low clipping
$L_{\mathrm{derate}}$ SOC ceiling $0.55$ $0.18$ Full throttle, SOC ceiling hit
The signal separation between $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ via $\delta_{\mathrm{throttle}}$ is |0.550.08|=0.47|0.55-0.08|=0.47, more than twice the combined standard deviation ($\sigma\approx 0.18$). This separation—identified in v1.5 as motivating a sixth observable—is the quantitative justification for elevating it to a state-level distinction in v2.
Circuit-dependent prior for $L_{\mathrm{derate}}$ at Melbourne.
At Melbourne (recharge ratio 1.0), the estimated fraction of $L$ laps that terminate in mandatory super-clipping is 0.40, giving a blended prior mean of $0.6\times 0.08+0.4\times 0.55=0.268$. In v2 this blended prior is eliminated: the HMM maintains separate posteriors over $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$, and Baum-Welch will estimate the circuit-specific transition rates between them directly from Race 1 telemetry.
@@ 4.3 Emission Signal Strength: Audit Results
ers level is strongly identifiable: $|\mu_{H}-\mu_{L_{\mathrm{harvest}}}|=6.0$ km/h in $\Delta v_{\mathrm{trap}}$ and 18 m in $\Delta b_{\mathrm{brake}}$, both substantially larger than the emission noise.
$L_{\mathrm{harvest}}$ vs. $L_{\mathrm{derate}}$ is identifiable via $\delta_{\mathrm{throttle}}$: separation of 0.47, more than twice the combined $\sigma$. This is the primary motivation for the v2 state-level split.
Tyre degradation signal is concentrated in the cliff state: $\Delta t_{\mathrm{sector}}\approx+0.8$ s and $\Delta b_{\mathrm{brake}}\approx+16$ m produce step changes the 5-lap baseline cannot absorb.
Override Mode status is not directly distinguishable from emissions in the non-active state (differences <0.5<0.5 km/h in $\Delta v_{\mathrm{trap}}$). The Override activation event is strongly identifiable ($\Delta v_{\mathrm{trap}}\approx+7$ km/h, $\Delta t_{\mathrm{sector}}\approx-0.65$ s).
@@ 4.4 Observation Discretisation
Observations are discretised to form a finite joint observation space as specified in Appendix A: $\Delta v_{\mathrm{trap}}$ (8 bins), $\Delta t_{\mathrm{sector}}$ (7 bins), $\Delta b_{\mathrm{brake}}$ (5 bins), $\sigma^{2}_{\mathrm{speed}}$ (5 bins), $z_{\mathrm{aero}}$ (2 bins), $\delta_{\mathrm{throttle}}$ (6 bins). Combined space: $8\times 7\times 5\times 5\times 2\times 6=16{,}800$ bins.
The emission matrix $E\in\mathbb{R}^{40\times 16800}$ is computed analytically by integrating the Gaussian, Bernoulli, and truncated-Gaussian distributions over bin boundaries, conditioning on each of the 40 hidden states. Two matrices are maintained: $E_{\mathrm{normal}}$ (outside activation zones) and $E_{\mathrm{zone}}$ (inside activation zones, with $\Delta v_{\mathrm{trap}}$ shifted by +4.0+4.0 km/h and $z_{\mathrm{aero}}$ non-zero).
@@ 5.1 Layer 1: The HMM Belief Update
The hmm has 40 hidden states, emission matrix $E\in\mathbb{R}^{40\times 16800}$, transition matrix $T\in\mathbb{R}^{40\times 40}$, and initial distribution $\pi\in\Delta^{40}$.
@@
The transition probability $T_{ij}=P(x_{t}[+1]=j\mid x_{t}=i)$ factors over the three components under conditional independence:
@@
The ers transition is marginalised over the rival’s action with estimated $P(\mathtt{burn})$:
$P(e^{\prime}\mid e)=p_{\mathrm{burn}}\cdot T^{\mathrm{burn}}_{ee^{\prime}}+(1-p_{\mathrm{burn}})\cdot T^{\mathrm{harvest}}_{ee^{\prime}}$ (14)
The $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ sub-modes have different transition dynamics (Appendix B): a car in $L_{\mathrm{harvest}}$ can transition to $M$ or $H$ under burn (it has reserve); a car in $L_{\mathrm{derate}}$ requires at least one sector of managed throttle to recover to $M$, and cannot transition directly to $H$. This asymmetry is strategically critical: it distinguishes a dangerous rival from a vulnerable one.
The Override Mode transition is proximity-conditioned. $P(m^{\prime}=\mathtt{available}\mid m=\mathtt{spent})=P(\mathrm{gap}<1\,\text{s})$, which is partially observable from live timing data. The tyre transition is monotone-absorbing. Full transition matrices are given in Appendix B.
@@
Phase 1 (pre-Melbourne): $E$ is set analytically; $T$ from physics-derived transition parameters; $\pi$ concentrates on $(e=H,m=\mathtt{available},\tau=\mathtt{new})$. Phase 2 (post-Melbourne): Baum-Welch EM (Baum et al., 1970) updates $E$ and $T$ from real race sequences.
@@ 5.2 Layer 2: The POMDP Policy
Input vector (66 dimensions).
Table 3: DQN input vector components. Total: 66 dimensions (increased from 55 in v1/v1.5 due to expansion of the belief vector from 30 to 40 states and expansion of own ers encoding from 3 to 4 modes).
@@
$e_{\mathrm{self}}$ one-hot 4 Own ERS level (4 modes in v2)
$m_{\mathrm{self}}$ one-hot 2 Own Override Mode status
$f_{\mathrm{self}}$ one-hot 3 Fuel (from lap number)
$\mathrm{gap\_ahead}$ one-hot 4 close / threat / neutral / clear
$\mathrm{gap\_behind}$ one-hot 4
$\delta_{\mathrm{gap}}$ scalar 1 Closure rate, normalised [1,1][-1,1]
$\mathrm{tyre\_compound}$ one-hot 3 soft / medium / hard
$\mathrm{tyre\_age}$ scalar 1 Normalised 0–1
$\mathrm{lap}$ scalar 1 Normalised 0–1
$\mathrm{sector}$ one-hot 3
$b_{t}$ probability vector 40 HMM belief state (40 states in v2)
Total 66
@@
$\text{Input}(66)\to\text{Dense}(256,\mathrm{ReLU})\to\text{Dense}(256,\mathrm{ReLU})\to\text{Dense}(128,\mathrm{ReLU})\to\text{Dense}(2)$
@@
Double DQN (van Hasselt et al., 2016) with experience replay (buffer capacity 50,000; Huber loss; gradient clipping at 10). $\varepsilon$-greedy exploration decays from 1.0 to 0.05 over synthetic pre-training.
@@
During synthetic pre-training only: $R_{\mathrm{shaped}}=R+\lambda\cdot\Phi$, $\lambda=0.3$, where
$\Phi=+0.10\cdot\mathbf{1}[\mathrm{MAP}(b_{t})=x^{*}_{t}]-0.20\cdot\mathbf{1}[\mathrm{trap\;condition}\wedge a=\mathtt{burn}]+0.05\cdot\mathbf{1}[\mathrm{gap\_ahead}\in\{\mathrm{close},\mathrm{threat}\}\wedge e_{\mathrm{self}}\geq M]$ (16)
@@
The dqn receives the hmm belief state $b_{t}$ rather than raw observables. In a pomdp, the belief state is a sufficient statistic of the entire observation history. In v2, the belief state now explicitly encodes $P(L_{\mathrm{harvest}})$ and $P(L_{\mathrm{derate}})$ as separate dimensions, directly actionable by the policy. The $\delta_{\mathrm{gap}}$ input provides a raw, hmm-independent signal for closure rate.
@@ 6.1 Data Source
Race telemetry is retrieved via the FastF1 Python library (Schaefer, 2020), which provides access to the official FOM timing and telemetry feed at 10 Hz sampling. The car_data endpoint provides per-car speed, throttle, and braking data at sector resolution.